# Edge-Spreading Raptor-Like LDPC Codes for 6G Wireless Systems

### v1

@@ VI-B Comparisons with 5G-NR LDPC Codes
Comparing the performance of ESRL codes and 5G-NR LDPC codes in a reasonably fair manner involves many factors. As stated in [3], computational complexity does not always accurately reflect the true silicon area cost, and throughput is significantly affected by the hardware architecture and the level of processing parallelism. Therefore, in addition to using the same code configurations, all decoding parameters and complexity should be the same in both software and hardware as much as possible.
To concentrate on the selection of channel coding without the influence of decoding algorithms and hardware architectures, we adopt the following approach. Specifically, we intentionally extend the uncoupled blocklength of ESRL codes to match that of 5G-NR LDPC codes. Combined with the multi-engine architecture in Section V-B, when SLME is configured as BP decoding on $\mathbf{H}_{\text{SC}}$, the corresponding ESRL decoder can be decomposed into $L$ sub-decoders operating in parallel (with the tail redundancy causing some extra overhead). Each sub-decoder is similar in size to that of a single 5G-NR LDPC decoder [22]. In this context, compared to multi-core 5G-NR LDPC decoders [50], we can run the same decoding algorithm with the same hardware implementation for ESRL codes to eliminate additional impacts.
@@
Figure 14: FER comparisons of ESRL codes and 5G-NR LDPC codes using L-SP decoding with $I_{\text{max}}=200$.
For high-throughput scenarios with $I_{\text{max}}=5$, as shown in Fig. 12, ESRL codes exhibit significant performance advantages compared to 5G-NR LDPC codes across various rates. For instance, at $R=0.8$, the ESRL code outperforms the 5G-NR LDPC code by $1.15$ dB, and although the gain slightly decreases it remains $0.65$ dB at $R=0.34$. Note that all algorithmic complexity and implementation complexity have been normalized, these gains are only attributable to channel coding. Moreover, to fully present the FER performance of 5G-NR LDPC codes, we refer to [49] and define such extended 5G-NR LDPC codes as $\mathcal{C}_{\text{5G}}^{\text{II}}$ $(K=84480)$ by redesigning a 5G-NR LDPC exponent matrix with a lifting size 10×10$\times$10 × larger $(Z=3840)$ and a non-decreasing girth. The FER of $\mathcal{C}_{\text{5G}}^{\text{II}}$, referred to as ‘5G-NR-Plus’, is plotted in Fig. 12. Given that the blocklength of the original 5G-NR LDPC codes is already on the order of tens of thousands, a 10×10$\times$10 × blocklength increase can bring only a $0.1$ dB gain compared to the original 5G-NR LDPC codes. Hence, ESRL codes still far outperform 5G-NR-Plus LDPC codes [49]. Subsequently, to intuitively highlight the difference between ESRL codes and previous protograph-based rate-compatible SC-LDPC code construction in [33, 34, 35], we replace the uncoupled $\mathbf{B}$ of ESRL codes with the 5G-NR matrix and define the code $\mathcal{C}_{\text{5G}}^{\text{III}}$. All other ESRL optimizations, such as the unified graph $\mathcal{G}_{\text{U}}$ and the tail matrix $\mathbf{Q}$, are still preserved in $\mathcal{C}_{\text{5G}}^{\text{III}}$ to ensure a fair comparison. This more classical SC-LDPC construction is referred to as ‘5G-NR-Coupled’ in Fig. 12. When taking $R=0.63$ as an example, edge spreading for the 5G-NR matrix brings a $0.4$ dB gain from coupling at the FER of $10^{-5}$, yet, compared to the ESRL code, the remaining $0.55$ dB gain is due to the well-designed ESRL code matrix. This demonstrates the significance of designing the optimal coupled matrix in ESRL codes, adhering to the design motivation behind ESRL codes.
Fig. 13 illustrates the relationship between $E_{b}/N_{0}$ at the FER of $10^{-5}$ against $I_{\text{max}}$ for ESRL codes and 5G-NR LDPC codes. At $R=0.52$, to match the ESRL performance of $I_{\text{max}}=5$, the 5G-NR LDPC code needs to enhance $I_{\text{max}}$ to $7$. This gap widens further when $R=0.8$. We summarize all comparisons of ESRL codes and 5G-NR LDPC codes in Table I. As shown in Table I, when $I_{\text{max}}=5$, the ESRL decoder can deliver a peak throughput of $209.4$ Gbps, while the 5G-NR LDPC decoder achieves a peak throughput of $195.4$ Gbps. These results are calculated by (15) of [22]. Notably, ESRL codes offer gains of more than $0.8$ dB at most rates. For 5G-NR with $I_{\text{max}}=7$, although the FER performance becomes comparable, the peak throughput drops to $139.5$ Gbps, indicating a $50.1\%$ throughput improvement for ESRL codes. For high-reliability scenarios, we use layered sum-product (L-SP) decoding [7] to replace L-NMS. Under $I_{\text{max}}=200$, ESRL codes outperform 5G-NR LDPC codes, especially at high code rates, e.g., this gain is $0.36$ dB at the FER of $10^{-5}$ when $R=0.87$. This greatly emphasizes the potential of ESRL codes for next-generation channel coding.
@@
    Peak throughout is attained with $R=0.87$.
@@ VII Conclusion
In this work, we propose a new SC-LDPC code family called ESRL LDPC codes, which concentrates on constructing an optimal matrix after coupling, resulting in outstanding performance while maintaining extensive rate compatibility. Compared to LDPC block codes, ESRL codes demonstrate superior performance, thanks to the convolutional structure of the coupled matrix, as well as high throughput potential due to the regular and compact structure of underlying block codes.
Beyond the structure, properties, and code profile of ESRL codes, we also present the complete design flow and optimization strategies, including a unified graph to optimize the edge-spreading process and a modified coupled RCA to analyze the decoding threshold. In addition, we propose a high-throughput decoding algorithm called SLME for ESRL codes, which can be decomposed into multiple sub-decoders running in parallel. It is noteworthy that we provide a design example of ESRL codes and conduct comprehensive comparisons with 5G-NR LDPC codes and their variants, such as extended or coupled 5G-NR LDPC codes. When all decoding parameters and complexity are the same, ESRL codes significantly outperform 5G-NR LDPC codes in terms of error rate and throughput, making them a promising solution for 6G wireless systems.

### v2

@@ VI-B Comparisons with 5G-NR LDPC Codes
Comparing the performance of ESRL codes and 5G-NR LDPC codes in a reasonably fair manner involves many factors. [^6] Our decoding framework for 5G-NR LDPC codes is available at [62]. As stated in [3], computational complexity does not always accurately reflect the true silicon area cost, and throughput is significantly affected by the hardware architecture and the level of processing parallelism. Therefore, in addition to using the same code configurations, all decoding parameters and complexity should be the same in both software and hardware as much as possible.
To concentrate on the selection of channel coding without the influence of decoding algorithms and hardware architectures, we adopt the following approach. Specifically, we intentionally extend the uncoupled blocklength of ESRL codes to match that of 5G-NR LDPC codes. Combined with the multi-engine architecture in Section V-C, when SLME is configured as BP decoding on $\mathbf{H}_{\text{SC}}$, the corresponding ESRL decoder can be decomposed into $L$ sub-decoders operating in parallel (with the tail redundancy causing some extra overhead). Each sub-decoder is similar in size to that of a single 5G-NR LDPC decoder [27]. In this context, compared to multi-core 5G-NR LDPC decoders [63], we can run the same decoding algorithm with the same hardware implementation for ESRL codes to eliminate additional impacts.
@@
For high-throughput scenarios with $I_{\text{max}}=5$, as shown in Fig. 13, ESRL codes exhibit significant performance advantages compared to 5G-NR LDPC codes across various rates. For instance, at $R=0.8$, the ESRL code outperforms the 5G-NR LDPC code by $1.15$ dB, and although the gain slightly decreases it remains $0.65$ dB at $R=0.34$. Although 5G-NR LDPC is designed to puncture two columns, we change its $\rho$ to $1$ to enrich the comparison. Under $I_{\text{max}}=5$, the decoding convergence of $\rho=1$ is better than that of $\rho=2$ in 5G-NR LDPC codes, especially at $R=0.87$. Yet, 5G-NR LDPC codes with $\rho=1$ are still worse than ESRL codes across various code rates. Note that all algorithmic complexity and implementation complexity have been normalized, these gains are only attributable to channel coding.
Moreover, to fully present the FER performance of 5G-NR LDPC codes, we refer to [61] and define such extended 5G-NR LDPC codes as $\mathcal{C}_{\text{5G}}^{\text{II}}$ $(K=84480)$ by redesigning a 5G-NR LDPC exponent matrix with a lifting size 10×10$\times$10 × larger $(Z=3840)$ and a non-decreasing girth. The FER of $\mathcal{C}_{\text{5G}}^{\text{II}}$, referred to as ‘5G-NR-Plus’, is plotted in Fig. 13. Given that the blocklength of the original 5G-NR LDPC codes is already on the order of tens of thousands, a 10×10$\times$10 × blocklength increase can bring only a $0.15$ dB gain compared to the original 5G-NR LDPC codes. ESRL codes still outperform 5G-NR-Plus LDPC codes [61]. Note that ESRL codes maintain better implementability than directly extending 5G-NR LDPC codes, since $Z=3840$ will bring higher routing complexity to the shift network in decoders for an extended code $\mathcal{C}_{\text{5G}}^{\text{II}}$. Subsequently, to intuitively highlight the difference between ESRL codes and previous protograph-based rate-compatible SC-LDPC code construction in [42, 43, 44], we replace the uncoupled $\mathbf{B}$ of ESRL codes with the 5G-NR matrix and define the code $\mathcal{C}_{\text{5G}}^{\text{III}}$. All other ESRL optimizations, such as the unified graph $\mathcal{G}_{\text{U}}$ and the tail matrix $\mathbf{Q}$, are still preserved in $\mathcal{C}_{\text{5G}}^{\text{III}}$ to ensure a fair comparison. This more classical SC-LDPC construction refers to ‘5G-NR-Coupled’ in Fig. 13. When taking $R=0.63$ as an example, edge spreading for the 5G-NR matrix brings a $0.4$ dB gain from coupling at the FER of $10^{-5}$. Even with one column punctured, compared to the ESRL code, there is still a gap of $0.45$ dB. This demonstrates the significance of designing the optimal coupled matrix in ESRL codes, adhering to the design motivation behind ESRL codes.
@@
Fig. 14 highlights FER comparisons of ESRL codes and 5G-NR LDPC codes with different $I_{\text{max}}$. Due to the design of 5G-NR LDPC codes intended for $20$ iterations, 5G-NR LDPC codes indeed perform better than ESRL codes in such cases. However, in other scenarios of both a limited ($I_{\text{max}}=5$) and a large number of iterations ($I_{\text{max}}=200$), ESRL codes show good performance over 5G-NR LDPC codes. Fig. 15 illustrates the relationship between $E_{b}/N_{0}$ at the FER of $10^{-5}$ against $I_{\text{max}}$ for ESRL codes and 5G-NR LDPC codes. At $R=0.52$, to match the ESRL performance of $I_{\text{max}}=5$, the 5G-NR LDPC code needs to enhance $I_{\text{max}}$ to $7$. This gap widens further when $R=0.8$. We summarize all comparisons of ESRL codes with 5G-NR LDPC codes in Table I. As shown in Table I, when $I_{\text{max}}=5$, the ESRL decoder can deliver a peak throughput of $209.4$ Gbps, while the 5G-NR LDPC decoder achieves a peak throughput of $195.4$ Gbps. These results are calculated by (15) of [27]. Notably, ESRL codes offer gains of more than $0.8$ dB at most rates. For 5G-NR with $I_{\text{max}}=7$, although the FER performance becomes comparable, the peak throughput drops to $139.5$ Gbps, indicating a $50.1\%$ throughput improvement for ESRL codes. For high-reliability scenarios, we use layered sum-product (L-SP) decoding [7] to replace L-NMS in Fig. 16. Under $I_{\text{max}}=200$, ESRL codes outperform 5G-NR LDPC codes, especially at high code rates, e.g., this gain is $0.36$ dB at the FER of $10^{-5}$ when $R=0.87$. This greatly emphasizes the potential of ESRL codes for next-generation channel coding.
Figure 16: FER comparisons of ESRL codes and 5G-NR LDPC codes [4] using L-SP decoding with $I_{\text{max}}=200$.
@@
    Peak throughput is attained with $R=0.87$.
@@
VI-C Performance of Windowed BP
ESRL codes can also support greater flexibility by windowed BP decoding [47]. Fig. 17 presents a comparison between windowed BP with various configurations and full BP. For brevity, let WBP-$(W,I_{\text{max}})$ denote windowed BP decoding with a window size $W$ and iteration count $I_{\text{max}}$. For instance, WBP-$(7,4)$ achieves a comparable computational complexity of full BP with $I_{\text{max}}=5$. While this setting underperforms full BP by $0.17$ dB at the FER of $10^{-5}$, it reduces hardware area by approximately $36.4\%$. Moreover, by increasing the iteration count from $I_{\text{max}}=4$ to $I_{\text{max}}=5$, the performance of a decoder with this window size is even better than full BP with $I_{\text{max}}=5$, due to the higher computational complexity of WBP-$(7,5)$. Note that since windowed BP utilizes less hardware and limits the message-passing to a smaller matrix dimension, its FER curve exhibits a slightly worse slope compared to full BP with the same number of iterations. With the same computational complexity, full BP delivers the best performance but comes with the highest hardware consumption. In contrast, windowed BP offers a broader design space exploration potential for ESRL codes, allowing optimization of parameters such as window size and sliding schedule to achieve a trade-off between performance and hardware complexity.
VI-D Combination with IR-HARQ
Let $\vartheta$ denote the system rate under IR-HARQ, i.e., defined as the average code rate of rate-compatible codes to satisfy a given frame error rate (FER) performance (decoding starts from a high code rate), where more details are available in (23) of [17]. This metric can indicate the number of practically transmitted coded bits in wireless systems in an IR-HARQ process. In combination with IR-HARQ, Fig. 18 shows that at a target FER of $10^{-5}$ using full BP decoding with $I_{\text{max}}=5$, the system rate $\vartheta$ of ESRL codes can be up to $31\%$ higher than that of 5G-NR LDPC codes at $E_{b}/N_{0}=3.0$ dB.
Figure 17: FER, normalized hardware area, and normalized computational complexity comparisons of windowed BP decoding and full BP decoding for ESRL codes with $R=0.8$ and $L=10$ under SLME decoding.
Figure 18: System rate comparison of ESRL codes and 5G-NR LDPC codes under IR-HARQ at a target FER of $10^{-5}$ with $I_{\text{max}}=5$.
@@ VII Conclusion
In this work, we propose an SC-LDPC code family called ESRL LDPC codes for 6G eMBB+ scenarios. The choice of an SC-LDPC code allows to extend and adjust the frame length by adjusting the coupling length in a flexible manner, without changing the matrix of the underlying block code. Larger block lengths thereby are not only in line with the desire to use larger transport block sizes in 6G eMBB+, but also enable high-throughput decoder architectures. To further enhance the throughput, the proposed ESRL code is optimized to not only achieve better error rate performance than 5G-NR LDPC codes with a high number of iterations, but also specifically with a low number of iterations. This property is achieved by focusing on the coupled matrix in the code construction, rather than the base matrix only from which the SC-LDPC code is derived. As an essential feature to be of interest to 6G eMBB+, ESRL codes maintain rate compatibility with a Raptor-like structure that naturally evolves from the 5G-NR LDPC code and that maintains compatibility with the IR-HARQ process that is essential to 6G eMBB+.
Further to the code construction and the proposal of a structure and code profile, we also present the complete design flow and optimization strategies, including a unified graph to optimize the edge spreading process and a modified coupled RCA to analyze the decoding threshold. To support the high-throughput requirement, we also propose a high-throughput decoding algorithm for ESLR codes, called SLME, which can be decomposed into multiple sub-decoders running in parallel to significantly enhance decoding parallelism. We provide a design example for a code and conduct comprehensive comparisons with 5G-NR LDPC codes and their variants, such as extended or coupled 5G-NR LDPC codes. When all decoding parameters and complexity are the same, the proposed code outperforms 5G-NR LDPC codes in terms of error rate and throughput in some specific scenarios (low and high number of iterations), which contributes a practical solution for applying SC-LDPC codes in future wireless systems.